# dMel: Speech Tokenization made Simple

### v2

@@ 4 Related Work
Despite these advancements, several challenges persist in the field of speech tokenization: i) balancing semantic and acoustic information in a unified representation; ii) reducing model complexity and computational requirements; iii) improving generalization to low-resource / out-of-domain data e.g. with mixed speech from multiple speakers, or multiple languages, or changing characteristics of recording equipment/sampling rate etc.; iv) developing truly single-stage tokenizers. Our proposed method, dMel, addresses these challenges by offering a training-free speech tokenization approach. By directly discretizing log mel-filterbanks into bins, it inherently preserves both semantic and acoustic information in a unified representation, while significantly reducing computational complexity.
@@
RichASR-RichTTS, joint model, 258M 7.6 20.0 4.4 1.9
@@ 5 Conclusion
In this work, we proposed dMel, a novel train-free speech tokenization method that discretizes log mel-filterbank energies directly into bins. By operating on the authentic log mel-filterbank representation, dMel inherently preserves both semantic and acoustic information in a unified tokenized representation. Our key contribution is the evaluation of dMel within a unified LM-style transformer architecture for speech recognition (ASR) and speech synthesis (TTS) tasks. Our dMel-based ASR model, RichASR, achieved the lowest word error rate among tokenization methods, robustly preserving semantic content. For TTS, dMel’s generation yielded the lowest WER, accurately reconstructing speech waveforms. Our dMel-based TTS model, RichTTS, achieved competitive naturalness, lowest error rates, and long audio generation capabilities.
@@ E.3 LM-Style Speech-to-Text
We found that span masking is key part of model training to enforce slef-attention to attend to speech part as well as to reduce exposure bias. The masking strategy is similar to the one used for TTS training: for every training step with probability $p$ the sample in the minibatch is masked with the mean span of 3 tokens with masking ration of 0.5. We found that the mean span of 1 token or 5 tokens gives the same results; while the mask probability $p$ is the most important hyper-parameter. The optimimal value for ASR is found to be 0.8, which is used in all final models.
@@ F.2 LM-Style Speech-to-Text
We noticd the results in [10] seems to be the SOTA for LM-style ASR model to the best of our knowledge. However, as many ablations are missed in [10], we took their open-sourced code and run ablations ourselves to have proper comparison with it. The final results, including ablation with dMel are shown in Table 15:

### v3

@@ 4 Related Work
Despite these advancements, several challenges persist in the field of speech tokenization: i) balancing semantic and acoustic information in a unified representation; ii) reducing model complexity and computational requirements; iii) improving generalization to low-resource / out-of-domain data e.g. with mixed speech from multiple speakers, or multiple languages, or changing characteristics of recording equipment/sampling rate etc.; iv) developing truly single-stage tokenizers. Our proposed method, dMel, addresses these challenges by offering a training-free speech tokenization approach. By directly discretizing log mel-filterbanks into bins, it inherently preserves both semantic and acoustic information in a unified representation, while significantly reducing computational complexity. Concurrently, [25] proposed a different mel-filterbanks based speech tokens: spectral codecs, where disjointed mel-bands are encoded separately and then quantized using an FSQ [29]. Although spectral codecs and dMel are both discretizing mel-filterbanks, dMel is encoder-free and tested with autoregressive generation tasks, while spectral codecs need encoders and is tested with non-autogressive generation task. Also, our scalar quantization method shares some similarities with FSQ, but FSQ is in a learned latent code space and necessitates an additional bound operation to limit the range of the latent codes.
@@
RichASR-RichTTS, joint model, 258M 7.5 15.3 4.4 1.9
@@ 5 Conclusion
In this work, we proposed dMel, a novel train-free speech tokenization method that discretizes log mel-filterbank energies directly into bins. By operating on the authentic log mel-filterbank representation, dMel inherently generalizes to out-of-domain data (e.g. speech with noise or other languages), preserves both semantic and acoustic information in a unified tokenized representation, and is streamable. Our key contribution is the evaluation of dMel within a unified LM-style transformer architecture for speech recognition (ASR) and speech synthesis (TTS) tasks. Our dMel-based ASR model, RichASR, achieved the lowest word error rate among tokenization methods, robustly preserving semantic content. For TTS, dMel’s generation yielded the lowest WER, accurately reconstructing speech waveforms. Our dMel-based TTS model, RichTTS, achieved competitive naturalness, lowest error rates, and long audio generation capabilities.
@@ E.3 LM-Style Speech-to-Text
We found that span masking is key part of model training to enforce self-attention to attend to speech part as well as to reduce exposure bias. The masking strategy is similar to the one used for TTS training: for every training step with probability $p$ the sample in the minibatch is masked with the mean span of 3 tokens with masking ration of 0.5. We found that the mean span of 1 token or 5 tokens gives the same results; while the mask probability $p$ is the most important hyper-parameter. The optimimal value for ASR is found to be 0.8, which is used in all final models.
@@ F.2 LM-Style Speech-to-Text
We noticed the results in [11] seems to be the SOTA for LM-style ASR model to the best of our knowledge. However, as many ablations are missed in [11], we took their open-sourced code and run ablations ourselves to have proper comparison with it. The final results, including ablation with dMel are shown in Table 14:
@@
Model Features dev-clean↓ dev-other↓
[16] (RNN-T – Conformer) Mel-10ms 1.9 4.1
[22] (CTC – Squeezeformer) Mel-10ms 2.3 5.8
Seq2Seq [14] Mel-10ms 2.4 5.4
dMel-10ms 2.5 5.9
Mel-25ms 2.8 6.5
dMel-25ms 2.7 6.2
CTC [15] Mel-10ms 2.1 5.4
dMel-10ms 2.1 5.6
Mel-25ms 2.1 5.4
dMel-25ms 2.3 6.1
LM-style dMel-25ms 3.4 9.5